OpenStudy (anonymous):

Ok. I need to write 2 equations. They both need a variable on each side of the equal sign. They also need to have either multiplication or division and addition or subtraction. The first equation needs to equal 8 when solved. The second equation needs to equal -2 when solved. Please help, I don't get it!

Parth (parthkohli):

The most basic form of both of these: \[\LARGE1) x+4 = 12\] \[\LARGE 2) x + 4 = 2\]

Parth (parthkohli):

Can I explain more or that's enough?

OpenStudy (anonymous):

I don't get that, actually. :T

Parth (parthkohli):

Do you know how to solve equations?

OpenStudy (anonymous):

The basic ones, yes. But were working on the equations with variables on both sides of the equal sign, and that just confuses me.

Parth (parthkohli):

Let me put variables on both sides, then do it.

OpenStudy (anonymous):

Oh wait, I see how you do the equations you posted before. I just don't know if my teacher'll take it because it doesn't have multiplication or division and addition or subtraction.

Parth (parthkohli):

Wait, it does. Look closely.

Parth (parthkohli):

\[\LARGE 1st => 2x - 2 = x + 6 \] \[\LARGE 2nd=> 2x - 2 = x - 4 \]

OpenStudy (anonymous):

Ok thank you, sorry I couldn't reply earlier, my keyboard wasn't working then I had to go to school.